OpenStudy (anonymous):

Write the equation of the circle whose center is at the origin and whose diameter is 12.

OpenStudy (solomonzelman):

center is at the origin, (that is center at point (0,0), so h=0, k=0) diameter is one half of the radius, so if diameter is 12 then radius is how long? (hope you got this part, and now I will remind you to equation of the circle form) \(\Large\color{black}{ \displaystyle (x-{\bf \color{red}{h}} )^2+ (y-{\bf \color{blue}{k}} )^2 = {\bf \color{green}{r}}^2 }\) Where 1) \(\large\color{black}{ \displaystyle({\rm \color{red}{h}} ,~{\rm \color{blue}{k}} )}\) is the \(\normalsize\color{red}{ \bf c\color{blue}{e}n\color{blue}{t}e\color{blue}{r} }\) of the circle. 2) \(\large\color{black}{ \displaystyle {\rm \color{green}{r}} }\) is the \(\normalsize\color{green}{ \bf radius }\) of the circle.

OpenStudy (anonymous):

There is no picture for me to find out what the middle of the circle is!?

OpenStudy (solomonzelman):

you don't need pics

OpenStudy (solomonzelman):

do you know how to fill out the \(\large\color{black}{ \displaystyle (x-{\rm \color{red}{h}} )^2+ (y-{\rm \color{blue}{k}} )^2 }\) ?

OpenStudy (anonymous):

no

OpenStudy (solomonzelman):

well, your center is (0,0) that means h=0 k=0

OpenStudy (solomonzelman):

something isn't clear?

OpenStudy (anonymous):

What do you mean!?

OpenStudy (solomonzelman):

I am asked you if you are confused about something in my explanation? I am asking you that because you don't seem to manage to fill out the left side of the equation of the circle ((( The equation of the circle: \(\large\color{black}{ \displaystyle (x-{\rm \color{red}{h}} )^2+ (y-{\rm \color{blue}{k}} )^2 = {\rm \color{green}{r}}^2 }\) )))

OpenStudy (anonymous):

Got it thanks!

OpenStudy (solomonzelman):

oh you got the equation of the circle? ok then...